An airplane flew for 3 hours with a tail wind of 40 km/hr. The return flight against the same wind took 4 hours. Find the speed of the airplane in still air.
Answer by Electrified_Levi(103)

Hi , Hope I can help
.
An airplane flew for 3 hours with a tail wind of 40 km/hr. The return flight against the same wind took 4 hours. Find the speed of the airplane in still air.
.
The airplane flew the same amount of miles each way.
Distance= (Rate)(Time)
.
We know the Time for each way, 3 hours with the wind, and 4 hours against the wind
.
We know that "x" is how fast it is going in still air.
.
Since the wind is adding to the speed on the first trip, we know that the speed is equal to (x+40)
.
On the return trip the plane is going against the wind, and decreasing the speed of the airplane, so the speed is equal to (x-40)
.
Now that we know both the time and the rate, we can solve it.
.
Distance = (Rate)(Time), or
(Rate)(Time)=(Rate)(Time)
.
Now we can put numbers in
.
(x + 40)(3) = (x - 40)(4)
3(x + 40) = 4(x - 40)
.
We will use distribution property,
.
3x + 120 = 4x - 160
.
We will now subtract 3x from both sides
.
0 + 120 = x - 160
.
We will add 160 to both sides
.
160 + 120 = x + 0
x = 280
.
The airplane is going 280 km/hr. in still air.
.
You can check by replacing "x" with "280"
.
3(x + 40) = 4(x - 40)
.
3(280 + 40) = 4(280 - 40)
3(320) = 4(240)
(960 = 960)
.
The airplane also went 960 miles, each way.
1,920 miles all together.
.
Hope I helped, Levi